# Bloom: Designing for LLM-Augmented Behavior Change Interactions

### v1

@@ Abstract.
Large language models (LLMs) offer novel opportunities to support health behavior change, yet existing work has narrowly focused on text-only interactions. Building on decades of HCI research demonstrating the effectiveness of UI-based interactions, we present Bloom, an application for physical activity promotion that integrates an LLM-based health coaching chatbot with established UI-based interactions. As part of Bloom’s development, we conducted a redteaming evaluation and contribute a safety benchmark dataset. In a four-week randomized field study (N=54) comparing Bloom to a non-LLM control, we observed important shifts in psychological outcomes: participants in the LLM condition reported stronger beliefs that activity was beneficial, greater enjoyment, and more self-compassion. Both conditions significantly increased physical activity levels, doubling the proportion of participants meeting recommended weekly guidelines, though we observed no significant differences between conditions. Instead, our findings suggest that LLMs may be more effective at shifting mindsets that precede longer-term behavior change.
@@ 1. Introduction
Recent advances in large language models (LLMs) (Ouyang et al., 2022; Bommasani et al., 2022) have sparked a growing interest in their applications for health behavior change, presenting novel opportunities for improved support. Early research on LLM-based health coaching chatbots suggest that they offer marked improvements in conversational flexibility over prior rule-based systems (Jörke et al., 2025; Wang et al., 2025; Xu et al., 2025; Heydari et al., 2025). This flexibility could allow LLMs to more accurately implement evidence-based health communication frameworks like motivational interviewing (Miller and Rollnick, 2023). Moreover, LLMs offer an improved capacity to interpret diverse sources of personal context, enabling interventions that are more effectively tailored to an individual (Khasentino et al., 2025; Merrill et al., 2024).
Unlike prior systems, which primarily rely on quantitative context (e.g., step count) for personalization, LLMs can also leverage qualitative context (Jörke et al., 2025) that is more easily expressed in natural language. Qualitative factors such as goals, motivation, values, life circumstances, time constraints, and access to resources are key constructs in behavior change theory and are crucial for effective personalization.
@@
Meanwhile, decades of research in human-computer interaction (HCI) has produced a rich body of interactions for health behavior change applications—from goal-setting (Ekhtiar et al., 2023) and self-tracking (Epstein et al., 2020) to ambient feedback (Consolvo et al., 2008; Murnane et al., 2023), and just-in-time nudges (Nahum-Shani et al., 2018)—which are both grounded in behavior change theory (Consolvo et al., 2009; Locke and Latham, 2002; Michie et al., 2011) and have demonstrated efficacy (Webb et al., 2010; Yang and Van Stee, 2019). However, existing research on LLMs for behavioral health has focused almost exclusively on text-only interactions, either through chat (Jörke et al., 2025; Wang et al., 2025; Xu et al., 2025), summaries (Merrill et al., 2024; Khasentino et al., 2025), or nudging (Song et al., 2025; Mantena et al., 2025). This narrow focus overlooks the potential for LLMs to augment established UI-based interactions, rather than replace them. Behavior change theory emphasizes that behavior is shaped not just by what is said, but when, how, and through which channels it is delivered (Mohr et al., 2014; Michie et al., 2011). HCI research has shown that multimodal interaction (Turk, 2014; Oviatt, 1999), which blends textual, graphical, ambient, or wearable delivered feedback, can help users manage cognitive load (Oviatt et al., 2004), make behaviors more salient through glanceable or ambient displays (Consolvo et al., 2008; Matthews, 2007; Gouveia et al., 2016), and foster affective engagement (Lin et al., 2006; Murnane et al., 2020). Taken together, this research suggests that behavior change interactions that augment LLM chat with other interaction modalities may offer greater efficacy.
@@
At the same time, the technology is recent and real-world evaluations of LLMs for behavioral health remain limited. Most studies have relied on static assessments or single-session studies (Mantena et al., 2025; Jörke et al., 2025; Meyer and Elsweiler, 2025; Srinivas et al., 2025; Heydari et al., 2025), while the few field trials conducted have included fewer than 20 participants (Wang et al., 2025; Xu et al., 2025). Importantly, most prior work lacks rigorous controls and assumes the efficacy of LLMs without directly evaluating LLM-based systems relative to established pre-LLM approaches. As a result, existing findings provide limited insights into the design requirements for LLM-augmented interactions or longitudinal LLM coaching.
@@
To address this gap, we present Bloom, an application for promoting PA that combines an LLM coaching chatbot with established UI-based behavior change interactions, including goal-setting, action planning, activity tracking, data visualization, an ambient display, and push notifications (Figure 1). Our goal in developing and evaluating Bloom was to surface design insights for LLM-augmented behavior change interactions and establish insights into their efficacy. Bloom’s LLM coach implements conversational strategies from motivational interviewing (Miller and Rollnick, 2023) and implements a scientifically validated health coaching program (King et al., 2002). Conversations with Bloom’s chatbot thus simultaneously serve as standalone interventions and as sources of qualitative context for other parts of the system to leverage for personalization. As part of Bloom’s development, we also conducted a redteaming evaluation with four domain experts and created a safety benchmark dataset for LLM coaching with 600 examples to evaluate our system’s safety filters.
@@
We evaluate Bloom in a four-week, between-subjects, randomized field study with $N=54$ participants, comparing Bloom to a no-LLM control that removes the LLM coach and all LLM augmentation. To the best of our knowledge, this is the first field study evaluation of LLM-based PA coaching. Our comparative study was formative and design-oriented, with the no-LLM control condition allowing us to isolate changes in behavior and experiences specifically attributable to LLM augmentation, as well as whether, when, and how LLMs provide benefits beyond well-designed, evidence-based, pre-LLM systems. Our evaluation uses a mixed-methods approach, including qualitative coding of semi-structured interviews; pre/post, weekly, and daily survey measures; objective PA data collected from participants’ wearables; as well as application usage logs including chats, plans, and UI interactions.
@@
We observed important shifts in psychological and motivational outcomes, with participants in the LLM condition reporting stronger beliefs that their activity was beneficial to their health. They also noted greater enjoyment of exercise, an expanded appreciation of what “counts” as activity, and increased self-compassion when goals were missed. The LLM condition also produced more varied and personalized plans, with slightly higher completion rates, and participants spent more than five times as much time in the app than control participants. Both conditions significantly increased their PA relative to a pre-study baseline, doubling the number of participants who met or exceeded the recommended 150 min/week of PA. However, we found no statistically significant differences in quantitative wearable PA outcomes between conditions during our four-week study period. Instead, our survey and interview findings suggest that LLMs may be particularly valuable for fostering positive mindsets, flexible planning, and sustained engagement—factors that behavior change theory links to longer-term maintenance rather than as immediate drivers of short-term levels of PA. We discuss the implications of these findings for designing LLM-augmented interventions and for understanding the role of LLMs in behavioral health systems.
@@
    The Bloom system, which implements a novel design for integrating LLM chat with established, UI-based behavior change interactions. Bloom is grounded in evidence-based strategies from behavior change theory and interactions from HCI research. Bloom leverages qualitative context from coaching conversations to personalize a number of LLM-augmented behavior change interactions.
@@
    The first safety evaluation of an LLM-based health coaching agent. We share results, prompts, and a benchmark dataset containing 600 examples to support future research on safety.
@@
    Findings from the first field study of an LLM-based PA coaching agent, a four-week, between-subjects randomized study with $N=54$ participants comparing Bloom to a non-LLM control. We analyze wearable PA outcomes, survey responses, system interaction logs, and qualitative coding of pre/post semi-structured interviews to assess both behavioral and psychological outcomes.
@@
    Design implications drawn from our field study’s findings to inform future work on LLM-augmented behavior change interactions and LLM health coaching systems.
@@ 2.1. HCI Systems for Physical Activity Promotion
Designing for PA promotion has a long history in HCI (Epstein et al., 2020), with many approaches advocated in the literature (Consolvo et al., 2014; Klasnja and Pratt, 2012; Hicks et al., 2023). One popular line of work draws from the broader field of personal informatics (Li et al., 2010) and aims to promote behavior change by helping users make sense of their activity patterns through self-monitoring (Korotitsch and Nelson-Gray, 1999; Choe et al., 2017) and reflection (Baumer, 2015; Bentvelzen et al., 2022; Cho et al., 2022). Common implementations frequently include visualizations (Aseniero et al., 2020; Choe et al., 2015; Kersten-van Dijk et al., 2017; Thudt et al., 2015) such as step count graphs, often seen in commercial health-tracking apps. Other systems employ more qualitative forms of feedback, including ambient displays (Consolvo et al., 2008; Murnane et al., 2023; Lane et al., 2012) and textual feedback (Bentley et al., 2013; Consolvo et al., 2006). Qualitative feedback can be less overwhelming (Daskalova et al., 2017) and promote more positive mindsets (Murnane et al., 2020) than quantitative feedback. Nudging represents another class of interventions, which in mobile health are often instantiated as push notification reminders for just-in-time support (Bentley and Tollmar, 2013; Intille, 2004; Vandelanotte et al., 2023; Nahum-Shani et al., 2018). Unlike other kinds of reflective or ambient support, just-in-time interventions aim to capture the user’s attention and often optimize intervention timing and content to maximize efficacy.
@@
Another prominent body of work focuses on goal-setting as a strategy for promoting PA (Ekhtiar et al., 2023; Munson and Consolvo, 2012). Goal-setting theory emphasizes that goals are most effective when they are personally meaningful, realistic, and sufficiently challenging (Locke and Latham, 2002). Users are often asked to set their own PA goals (Gasser et al., 2006; Lin et al., 2006), or are alternatively provided a goal based on their behavior history (Consolvo et al., 2006). Prior work has also employed conversational agents for collaborative goal-setting (King et al., 2020; Bickmore et al., 2005; Jörke et al., 2025). Since intentions alone do not always translate to goal achievement, action plans (Agapie et al., 2018) and implementation intentions (Gollwitzer and Sheeran, 2006; Gollwitzer, 1999; Sefidgar et al., 2024) are popular techniques for facilitating goal realization by specifying when, where, and how a behavior will occur.
@@
Beyond individual-level interventions, prior work has also explored social strategies for behavior change. These strategies typically involve sharing within a social network (Consolvo et al., 2006; Lin et al., 2006; Toscos et al., 2006; Anderson et al., 2007) or incorporating leaderboards and challenges to introduce competition in games (Buis et al., 2009; Göbel et al., 2010; Sinclair et al., 2007; Althoff et al., 2016; Shameli et al., 2017). While collaborative approaches have been shown to improve engagement and activity levels in some contexts, social comparison can sometimes feel demotivating or misaligned with preferences (Munson and Consolvo, 2012; Murnane et al., 2015), making social features challenging to design effectively. Competitive elements in particular, such as leaderboards or challenges, have shown mixed effects on users’ motivation, especially during periods when they were less active (Klasnja et al., 2009; Consolvo et al., 2006). Although gamification can increase short-term engagement, its long-term impact on sustained behavior change also remains underexplored (Stepanovic and Mettler, 2018).
@@
Bloom implements several interactions from this literature, including activity tracking, data visualizations, ambient displays, push notifications, goal setting, and action planning. Due to the challenge of appropriately designing for social and gamification interactions, compounded with the uncertainty of designing for novel, LLM-augmented interactions, we exclude these interactions, leaving them as important explorations for future work.
@@ 2.2. Health Coaching
Health coaching is an established intervention for facilitating behavior change (Olsen and Nesbitt, 2010; Wolever et al., 2013; Olsen, 2014). Although coaching is a multi-disciplinary concept with numerous everyday associations (e.g., sports, executive, or life coaching), here we adopt Olsen et al.’s (Olsen, 2014) definition of health coaching as “a goal-oriented, client-centered partnership that is health-focused and occurs through a process of client enlightenment and empowerment.” A common framework used in such programs is motivational interviewing (MI) (Miller and Rollnick, 2023), an evidence-based counseling approach that emphasizes eliciting motivation from the client rather than imposing prescriptive advice.
@@
However, in-person health coaching is expensive, not widely accessible, and does not scale to global need. A long history of research on health dialogue systems (Bickmore and Giorgino, 2006) and automated health coaching (Mitchell et al., 2021) has aimed to address these challenges by emulating aspects of human coaching with conversational agents. This research dates back at least as far as the seminal work on the Transtheoretical Model (TTM) itself (Prochaska and Velicer, 1997), and we refer to (Jörke et al., 2025; Mitchell et al., 2021) for reviews of pre-LLM automated coaching. While meta-analyses have found that pre-LLM coaching chatbots are effective at promoting physical activity (Singh et al., 2023), these systems rely on predefined rules or templates that inherently limit their flexibility and personalization compared to human coaches (Luo et al., 2021). Unlike prior rule-based systems, Bloom leverages an LLM to engage in dynamic, open-ended conversations, interpret and integrate user data through tool calls, and reference past interactions.
2.3. Large Language Models and Behavioral Health

### v2

@@ Abstract.
Large language models (LLMs) offer novel opportunities to support health behavior change, yet existing work has narrowly focused on text-only interactions. Building on decades of HCI research on effective behavior change interactions, we present Bloom, an application for physical activity promotion that integrates an LLM-based health coaching chatbot with existing design strategies and UI elements. As part of Bloom’s development, we conducted a redteaming evaluation and contribute a safety benchmark dataset. In a four-week randomized field study (N=54) comparing Bloom to a no-LLM control, we observed important shifts in psychological outcomes: participants in the LLM condition reported stronger beliefs that activity was beneficial, greater enjoyment, and more self-compassion. Both conditions significantly increased physical activity levels, doubling the proportion of participants meeting recommended weekly guidelines, though descriptively, we observed no advantage for the LLM condition in short-term physical activity levels. Instead, our findings suggest that LLMs may be more effective at shifting mindsets that precede longer-term behavior change.
Physical activity, health coaching, behavior change, large language models (LLM), mobile health
copyright: noneconference: the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spaindoi: 10.1145/3772318.3790506isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Natural language interfacesccs: Applied computing Health informaticsccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Computing methodologies Natural language processing
Figure 1. Bloom integrates an LLM-based physical activity coaching chatbot (Beebo) with established behavior change interactions. (A) An onboarding conversation with Beebo, an LLM health coaching chatbot. (B) The app’s home screen, showing weekly progress, upcoming activities, and proactive messages from Beebo. (C) LLM-generated activity plans and progress summaries. (D) LLM-generated notifications encouraging action or reflection. (E) A garden-based ambient display on the user’s lockscreen.
Overview of the Bloom application. A (left): A conversation between the user and Beebo, the LLM health coach, depicted as a cartoon bee. The user shares that they recently had a baby and expresses a desire for more energy to play. Beebo responds by affirming their motivations and asks what kinds of activities they would like to be able to do in the future. B (center): The homescreen (“Today” screen) of the application, showing a garden ambient display with widgets for weekly progress, upcoming workouts, and a proactive message from Beebo. C (top right): A plan widget, showing a calendar with Monday/Wednesday/Friday walks and a Saturday bike ride, and a progress summary where Bloom congratulates the user for completing three of four workouts. D (bottom right): A push notification reminding the user about their scheduled family bike ride. E (far right): A garden-based ambient display on an iPhone lockscreen, showing a prominent garden wallpaper as well as weekly progress and the next workout widgets.
@@ 1. Introduction
Recent advances in large language models (LLMs) (Ouyang et al., 2022; Bommasani et al., 2022) have sparked a growing interest in their applications for health behavior change, presenting novel opportunities for improved support. Early research on LLM-based health coaching chatbots suggests that they offer marked improvements in conversational flexibility over prior rule-based systems (Jörke et al., 2025; Wang et al., 2025; Xu et al., 2025; Heydari et al., 2025). This flexibility could allow LLMs to more accurately implement evidence-based health communication frameworks like motivational interviewing (Miller and Rollnick, 2023). Moreover, LLMs offer an improved capacity to interpret diverse sources of personal context, enabling interventions that are more effectively tailored to an individual (Khasentino et al., 2025; Merrill et al., 2026). Unlike prior systems, which primarily rely on quantitative context (e.g., step count) for personalization, LLMs can also leverage qualitative context (Jörke et al., 2025) that is more easily expressed in natural language. Qualitative factors such as goals, motivation, values, life circumstances, time constraints, and access to resources are key constructs in behavior change theory and are crucial for effective personalization.
@@
Meanwhile, decades of research in human-computer interaction (HCI) has produced a rich body of interactions for health behavior change applications—from goal setting (Ekhtiar et al., 2023) and self-tracking (Epstein et al., 2020) to ambient feedback (Consolvo et al., 2008; Murnane et al., 2023), and just-in-time nudges (Nahum-Shani et al., 2018)—which are both grounded in behavior change theory (Consolvo et al., 2009; Locke and Latham, 2002; Michie et al., 2011) and have demonstrated efficacy (Webb et al., 2010; Yang and Van Stee, 2019). However, existing research on LLMs for behavioral health has focused almost exclusively on text-only interactions, either through chat (Jörke et al., 2025; Wang et al., 2025; Xu et al., 2025), summaries (Merrill et al., 2026; Khasentino et al., 2025), or nudging (Song et al., 2025; Mantena et al., 2025). This narrow focus overlooks the potential for LLMs to augment established interactions, rather than replace them. Behavior change theory emphasizes that behavior is shaped not just by what is said, but when, how, and through which channels it is delivered (Mohr et al., 2014; Michie et al., 2011). HCI research has shown that multimodal interaction (Turk, 2014; Oviatt, 1999), which blends textual, graphical, ambient, or wearable delivered feedback, can help users manage cognitive load (Oviatt et al., 2004), make behaviors more salient through glanceable or ambient displays (Consolvo et al., 2008; Matthews, 2006; Gouveia et al., 2016), and foster affective engagement (Lin et al., 2006; Murnane et al., 2020). Taken together, this research suggests that behavior change interactions that augment LLM chat with other interaction modalities may offer greater efficacy.
@@
At the same time, the technology is recent and real-world evaluations of LLMs for behavioral health remain limited. Most studies have relied on static assessments or single-session studies (Mantena et al., 2025; Jörke et al., 2025; Meyer and Elsweiler, 2025; Srinivas et al., 2025; Heydari et al., 2025), with few studies reporting on preliminary field deployments (Wang et al., 2025; Xu et al., 2025; Loerakker et al., 2025), offering only partial insights into how users engage with LLM-based behavior change systems over time in real-world settings. Importantly, most prior work assumes the efficacy of LLMs without directly evaluating LLM-based systems relative to established pre-LLM approaches. As a result, this limits our understanding of the unique design opportunities and requirements for LLM-augmented behavior change systems.
@@
To address this gap, we present Bloom, a mobile application for promoting PA that combines an LLM coaching chatbot with established behavior change interactions, including goal setting, action planning, activity tracking, data visualization, an ambient display, and push notifications (Figure 1). Our goal in developing and evaluating Bloom was to surface design insights for LLM-augmented behavior change interactions by studying how LLM augmentation shapes user experience and behavior. Bloom’s LLM coach implements conversational strategies from motivational interviewing (Miller and Rollnick, 2023) and a scientifically validated health coaching program (King et al., 2002). Conversations with Bloom’s chatbot simultaneously serve as standalone interventions and sources of qualitative context for other parts of the system to use for personalization. As part of Bloom’s development, we also conducted a redteaming evaluation with four domain experts and created a safety benchmark dataset for LLM coaching with 600 examples to evaluate our system’s safety filters.
@@
We evaluate Bloom in a four-week, between-subjects, randomized field study with $N=54$ participants, comparing Bloom to a no-LLM control that removes the LLM coach and all LLM augmentation. Our comparative study was formative, exploratory, and primarily design-oriented, with the no-LLM control condition allowing us to isolate changes in participant experiences specifically attributable to LLM augmentation. A secondary aim was to gather preliminary evidence of Bloom’s impact on measurable PA outcomes, noting that our early-stage study was not designed to establish statistically significant treatment-control differences (e.g., as a clinical feasibility or efficacy trial (Czajkowski et al., 2015)). Our evaluation uses a mixed-methods approach, including qualitative coding of semi-structured interviews; pre/post, weekly, and daily survey measures; objective PA data collected from participants’ wearables; and application usage logs including chats, plans, and UI interactions.
@@
We observed important shifts in psychological and motivational outcomes, with participants in the LLM condition reporting stronger beliefs that their activity was beneficial to their health. They also noted greater enjoyment of exercise, an expanded appreciation of what “counts” as activity, and increased self-compassion when goals were missed. The LLM condition also produced more varied and personalized plans, with slightly higher completion rates, and participants spent more than five times as much time in the app than control participants. Both conditions significantly increased their PA relative to a pre-study baseline, doubling the number of participants who met or exceeded the recommended 150 min/week of PA. However, across quantitative wearable outcomes, descriptive patterns did not indicate a clear advantage of the LLM condition during the four-week study period. Instead, our survey and interview findings suggest that LLMs may be particularly valuable for fostering positive mindsets, flexible planning, and sustained engagement—factors that behavior change theory links to longer-term maintenance rather than as immediate drivers of short-term levels of PA. We discuss the implications of these findings for designing LLM-augmented interventions, highlighting how the LLM’s role as a conversational coach shifted mindsets and motivation more than short-term PA, how qualitative context enabled non-prescriptive interactions that promote agency, and how relational cues increased engagement while introducing design trade-offs around attachment and overreliance.
@@
    The Bloom system, which implements a novel design for integrating an LLM health coaching chatbot with established design strategies and interface elements. Bloom is grounded in evidence-based strategies from behavior change theory and interactions from HCI research. Bloom leverages qualitative context from coaching conversations to personalize a number of LLM-augmented behavior change interactions.
@@
    A safety evaluation of our LLM-based health coaching agent. We share results, prompts, and a benchmark dataset containing 600 examples to support future research on safety.
@@
    Findings from our field study of an LLM-based PA coaching agent, a four-week, between-subjects randomized study with $N=54$ participants comparing Bloom to a no-LLM control. We analyze qualitative coding of pre/post semi-structured interviews, survey responses, wearable PA outcomes, and system interaction logs to assess both behavioral and psychological outcomes.
@@
    Design implications drawn from our field study’s findings to inform future work on LLM-augmented behavior change systems and LLM health coaching.
@@ 2.1. HCI Systems for Physical Activity Promotion
Designing for PA promotion has a long history in HCI (Epstein et al., 2020), with many approaches advocated for in the literature (Consolvo et al., 2014; Klasnja and Pratt, 2012; Hicks et al., 2023). One popular line of work draws from the broader field of personal informatics (Li et al., 2010) and aims to promote behavior change by helping users make sense of their activity patterns through self-monitoring (Korotitsch and Nelson-Gray, 1999; Choe et al., 2017) and reflection (Baumer, 2015; Bentvelzen et al., 2022; Cho et al., 2022). Common implementations frequently include visualizations (Aseniero et al., 2020; Choe et al., 2015; Kersten-van Dijk et al., 2017; Thudt et al., 2015) such as step count graphs, often seen in commercial health tracking apps. Other systems employ more qualitative forms of feedback, including ambient displays (Consolvo et al., 2008; Murnane et al., 2023; Lane et al., 2012) and textual feedback (Bentley et al., 2013; Consolvo et al., 2006). Qualitative feedback can be less overwhelming (Daskalova et al., 2017) and promote more positive mindsets (Murnane et al., 2020) than quantitative feedback. Nudging represents another class of interventions, which in mobile health are often instantiated as push notification reminders for just-in-time support (Bentley and Tollmar, 2013; Intille, 2004; Vandelanotte et al., 2023; Nahum-Shani et al., 2018). Unlike other kinds of reflective or ambient support, just-in-time interventions aim to capture the user’s attention and often optimize intervention timing and content to maximize efficacy.
@@
Another prominent body of work focuses on goal setting as a strategy for promoting PA (Ekhtiar et al., 2023; Munson and Consolvo, 2012). Goal setting theory emphasizes that goals are most effective when they are personally meaningful, realistic, and sufficiently challenging (Locke and Latham, 2002). Users are often asked to set their own PA goals (Gasser et al., 2006; Lin et al., 2006), or are alternatively provided a goal based on their behavior history (Consolvo et al., 2006). Prior work has also employed conversational agents for collaborative goal setting (King et al., 2020; Bickmore et al., 2005; Jörke et al., 2025). Since intentions alone do not always translate to goal achievement, action plans (Agapie et al., 2018) and implementation intentions (Gollwitzer and Sheeran, 2006; Gollwitzer, 1999; Sefidgar et al., 2024) are popular techniques for facilitating goal realization by specifying when, where, and how a behavior will occur.
@@
Beyond individual-level interventions, prior work has also explored social strategies for behavior change (Saksono and Parker, 2024). These strategies typically involve sharing within a social network (Consolvo et al., 2006; Lin et al., 2006; Toscos et al., 2006; Anderson et al., 2007) or incorporating leaderboards and challenges to introduce competition in games (Buis et al., 2009; Göbel et al., 2010; Sinclair et al., 2007; Althoff et al., 2016; Shameli et al., 2017). While collaborative approaches have been shown to be effective, social comparison can sometimes feel demotivating or misaligned with preferences (Munson and Consolvo, 2012; Murnane et al., 2015), making social features more challenging to design effectively. Competitive elements in particular, such as leaderboards or challenges, have shown mixed effects on users’ motivation, especially during periods when they were less active (Klasnja et al., 2009; Consolvo et al., 2006). Although gamification can increase short-term engagement, its long-term impact on sustained behavior change also remains underexplored (Stepanovic and Mettler, 2018).
@@
Bloom implements several behavior change interactions from this literature, including activity tracking, data visualizations, ambient displays, push notifications, goal setting, and action planning. Due to the challenge of appropriately designing for social and gamification interactions, compounded with the uncertainty of designing for novel, LLM-augmented interactions, we exclude these interactions as important explorations for future work.
@@ 2.2. Health Coaching
Health coaching is an established intervention for facilitating behavior change (Olsen and Nesbitt, 2010; Wolever et al., 2013; Olsen, 2014). Although coaching is a multidisciplinary concept with numerous everyday associations (e.g., sports, executive, or life coaching), here we adopt Olsen et al.’s (Olsen, 2014) definition of health coaching as “a goal-oriented, client-centered partnership that is health-focused and occurs through a process of client enlightenment and empowerment.” A common framework used in such programs is motivational interviewing (Miller and Rollnick, 2023), an evidence-based counseling approach that emphasizes eliciting motivation from the client rather than imposing prescriptive advice.
@@
However, in-person health coaching is expensive, not widely accessible, and does not scale to global need. A long history of research on health dialogue systems (Bickmore and Giorgino, 2006) and automated health coaching (Mitchell et al., 2021) has aimed to address these challenges by emulating aspects of human coaching with conversational agents. This research dates back at least as far as the seminal work on the Transtheoretical Model (TTM) itself (Prochaska and Velicer, 1997), and we refer to (Jörke et al., 2025; Mitchell et al., 2021) for reviews of pre-LLM automated coaching. While meta-analyses have found that pre-LLM coaching chatbots are effective at promoting PA (Singh et al., 2023), these systems rely on predefined rules or templates that inherently limit their flexibility and personalization compared to human coaches (Luo et al., 2021). Unlike prior rule-based systems, Bloom leverages an LLM to engage in dynamic, open-ended conversations, interpret and integrate user data through tool calls, and reference past interactions.
2.3. LLMs and Behavioral Health